OpenStudy (anonymous):

whats -3x=+y=4 in slope intercept form?

OpenStudy (anonymous):

just use simple algebra to get your y on one side and your x on another, then if you want to get rid of that negative sign multiple BOTH sides by -1

OpenStudy (anonymous):

Are you sure you typed the problem correctly?

OpenStudy (anonymous):

-3+y=4 sorry

jigglypuff314 (jigglypuff314):

if you mean -3x + y = 4 to put it into slope-intercept form you solve for y in other words you try to get y all alone on one side of the equals sign so for -3x + y = 4 you would add 3x to both sides +3x +3x so y = 3x + 4

OpenStudy (anonymous):

what you need to do is simple, slope-intercept form is y=mx+b, so simply add or subtract until you get the values you want on their side

OpenStudy (anonymous):

thanks